OpenStudy (jbeanz97):

can someone please help me with an algebra1 question?

OpenStudy (jbeanz97):

\[if, f(x)=\frac{ 2x }{ 2x+1 } and, g(x)=\frac{ 1 }{ 2x+1 }, find f(x)+g(x)\]

geerky42 (geerky42):

What did you get?

OpenStudy (jbeanz97):

i dont understand how to solve it

geerky42 (geerky42):

You just need to find and simplify f(x)+g(x)

geerky42 (geerky42):

Do you know what f(x) and g(x) are?

OpenStudy (jbeanz97):

well it says f(x) and g(x) are those fractions, but how do i simplify those fractions?

geerky42 (geerky42):

Well, you have \(\dfrac{2x}{2x+1}+\dfrac{1}{2x+1}\). Do you know what to do next? Note that denominators are same

OpenStudy (jbeanz97):

do we cross cancel?

geerky42 (geerky42):

Say you have like \(\dfrac{2}{5}+\dfrac{3}{5}\) to solve, what will you do?

OpenStudy (jbeanz97):

add the top and bottom?

OpenStudy (jbeanz97):

so it would give you 4/10?

OpenStudy (jbeanz97):

**5/10

geerky42 (geerky42):

no, it would give you 5/5 when denominator is same, you can add numerators, only numerators.

OpenStudy (jbeanz97):

oohh okay i gotcha

geerky42 (geerky42):

So adding numerator together and you will have \(\dfrac{2x+1}{2x+1}\). make sense?

OpenStudy (jbeanz97):

yeah!

geerky42 (geerky42):

so what's next?

OpenStudy (jbeanz97):

hmm im not sure :/

geerky42 (geerky42):

you do noticed that numerator and denominator are same?

OpenStudy (jbeanz97):

yeah

OpenStudy (jbeanz97):

so would it become 1

geerky42 (geerky42):

right! so we have \(\dfrac{\cancel{2x+1}}{\cancel{2x+1}} = \dfrac{1}{1}=\boxed{1}\)

geerky42 (geerky42):

So we have f(x)+g(x)=1 that's your answer. is that clear?

OpenStudy (jbeanz97):

yes that was clear! thanks for the help!! :)

geerky42 (geerky42):

welcome

OpenStudy (jbeanz97):

could you possibly help me with another question?